OpenStudy (anonymous):

A rectangle is three times as long as it is wide. If the length is decreased by 2 cm and the width is increased by 3 cm, the new rectangle formed has an area of 42 cm2. What are the dimensions of the original rectangle? A. 9 cm × 3 cm B. 7 cm × 6 cm C. 21 cm × 7 cm D. 15 cm × 5 cm

OpenStudy (anonymous):

Can anybody please help?

OpenStudy (mathmate):

There are two ways to approach this problem. The first way is to set up a system of (2) equations for the length and the width, and solve accordingly. The second way takes advantage of the given answers. You only have to give 4 tries according to instructions. First eliminate answers where length is not 3 times the width. That leaves us with three answers to check. take the first one: L=9, W=3 Reduce L by 2 and increase W by three to get 7*6=42 bingo, that satisfy the conditions. If you want, you can check the rest.

OpenStudy (anonymous):

Is C the correct answer? @mathmate

OpenStudy (mathmate):

For C: L=21, W=7 The questions says (L-2)*(W+3)=42 Let's check: (21-2)*(7+3)=19*10=190 does not equal 42, so it is not the right answer. For your benefit, would you kindly reread my previous post.

OpenStudy (anonymous):

it is A!!!

OpenStudy (mathmate):

Exactly! Good job! :)

OpenStudy (anonymous):

(L-2)x(W+3) = (9-2) x (3+3) = 42

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mathmate):

yw!